OpenStudy (anonymous):

The volume of the solid obtained by rotating the region bounded by y=x^2, \ y = 6 x about the line y = 36 can be computed using the method of washers or disks via an integral

OpenStudy (anonymous):

Have you done it? You'll have to find the radius of the washer..

OpenStudy (anonymous):

yeah...well i know the limit of integration it is 0 to 6

OpenStudy (anonymous):

i'm trying to find the V

OpenStudy (anonymous):

not in figure but just how to set it up

OpenStudy (anonymous):

this i my answer so far pi [(6 - (1/6)y)^2-(6-sqrt(y))^2]

OpenStudy (anonymous):

but not sure

OpenStudy (anonymous):

the answer ask for that kind of set up :)

OpenStudy (anonymous):

my answer is slightly different though. Why do you take dy instead of dx considering that it is rotating about horizontally?

OpenStudy (anonymous):

sorry it dx...

OpenStudy (anonymous):

you're right

OpenStudy (anonymous):

It simplifies the equation alot :)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what's ur answer?

OpenStudy (anonymous):

I'm getting something like \(V=\pi \int\limits_{0}^{6}( 72+6x-x^2)dx\)

OpenStudy (dumbcow):

hmm... \[V = \pi \int\limits_{0}^{6}(36-x^{2})^{2} - (36-6x)^{2} dx\]